Not logged in
Thu Feb 12 17:43:33 2009
SRC strangeness solved
I found the problem with the DRMI lock I had last night was caused by the zero gain in the PD11_I filter.
I don't know how it happened but putting it back to 1.000 made the DRMI lock far more stable and AS166Q got more than 3000.
I also re-centered POY PD to remove the offset in the y-arm loop. The large power drops while y-arm is locked by itself were eliminated.